Could you please help me out?
A square and an equilateral triangle have the same perimeter. Each side of the triangle is 4 inches longer than each side of the square. What is the perimeter of the figures? What is the area of the square?

Answer by ankor@dixie-net.com(22740)

A square and an equilateral triangle have the same perimeter. Each side of the triangle is 4 inches longer than each side of the square. What is the perimeter of the figures?
:
Let x = side of the square
then
(x+4) = side of the triangle
;
Perimeter of the square = 4x
Perimeter of the triangle = 3(x+4)
:
It says they are equal
4x = 3(x+4)
4x = 3x + 12
4x - 3x = 12
x = 12 is the side of the square (4*12=48)
and
12 + 4 = 16 is the side of the triangle (3*16=48)
;
What is the area of the square?
12^2 = 144 inches